# The action of an inverse semigroup on its Stone-Čech compactification

### v1

@@ Abstract.
We initiate the study of the Stone-Čech transformation groupoid $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ of an inverse semigroup $\mathcal{S}$. This groupoid occurs as a model for the inverse semigroup Roe algebra $\mathcal{R}_{\mathcal{S}}$ of Lledo and Martinez; in particular, the action of $\mathcal{S}$ on $\beta\mathcal{S}$ is amenable in the sense of Exel and Starling if and only if the reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact. We prove that the properties of being Hausdorff, principal, and effective are all equivalent for $\mathcal{G}$, and give an algebraic condition on $\mathcal{S}$ equivalent to the Hausdorffness of $\mathcal{G}$. Finally, we show that the Hausdorffness of Exel’s tight groupoid $\mathcal{G}_{\text{tight}}(\mathcal{S})$ is necessary for the Hausdorffness of $\mathcal{G}$.
@@ 1. Introduction
A semigroup $\mathcal{S}$ is called inverse if, for every $s\in\mathcal{S}$, there is a unique $s^{*}\in\mathcal{S}$ with $ss^{*}s=s$ and $s^{*}ss^{*}=s^{*}$. The dynamics and geometry of inverse semigroups have been productively studied in a way that generalizes the discrete group setting. In [lledomartinez], Lledo and Martinez described the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$ of a countable inverse semigroup $\mathcal{S}$, and showed that this C-algebra was naturally isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
In this note, we describe an étale groupoid model for $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. Like the group case, this étale groupoid arises from a left-translation action of $\mathcal{S}$ on its Stone-Čech compactification (Theorem 3.1). It follows immediately that Exel’s reduced inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable as an étale groupoid. Said differently, the following generalization of the discrete group case ([ozawa], Theorem 3) holds:
Theorem 1.
Exel’s inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable in the sense of Exel and Starling [starling].
It is well-known that the action of a discrete group $\Gamma$ on $\beta\Gamma$ is free (see for example [NR14, Example 2.14]. In the inverse semigroup setting, one may therefore anticipate that $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ is always principal. Using a finite cover criterion of Exel and Pardo [EP16], we obtain the following generalization of this result.
@@
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff,
@@
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal,
@@
This generalizes the group case because $\mathcal{S}\ltimes\beta\mathcal{S}$ is always Hausdorff when $\mathcal{S}$ is a group. Using Theorem 2, we give a criterion on $\mathcal{S}$ (Theorem 3.8) which implies $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff for the class of abstract pseudogroups (Corollary 3.9). Finally, since there are many examples of inverse semigroups with non-Hausdorff tight groupoids, condition (3) of Theorem 2 provides many examples of inverse semigroups for which $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff nor principal.
@@ 2. Preliminaries
In what follows, $\mathcal{S}$ denotes a countable inverse semigroup. An element $e\in\mathcal{S}$ is called idempotent if $e^{2}=e$. For any $\in\mathcal{S}, ss$$^{*}s$ and $ss^{*}$ are idempotent. Denote by $E_{\mathcal{S}}$ the set of idempotent elements of $\mathcal{S}$. It is a pleasant fact that $E_{\mathcal{S}}$ is a commutative subsemigroup of $\mathcal{S}$, and that $\mathcal{S}$ is a group if and only if $E_{\mathcal{S}}$ is a singleton. If $X$ is a set, then the collection of all partial bijections between subsets of $X$, denoted by $\mathcal{I}(X)$, is an inverse semigroup under the operation of partial composition called the symmetric inverse monoid on $X$.
@@
For any poset $(P,\prec)$ and $A\subseteq P$ we will use the notation
@@
We now introduce actions of inverse semigroups on topological spaces and C-algebras. Although this note is only interested in two such actions (one on $\beta\mathcal{S}$, and one on $\ell^{\infty}(\mathcal{S})$), we feel it will be beneficial to introduce the theory in somewhat excessive generality. Let $\mathcal{A}$ be a C-algebra. An action of an inverse semigroup $\mathcal{S}$ on $\mathcal{A}$ is a semigroup homomorphism $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{A})$ such that, for each $s\in\mathcal{S}$,
  1. (1)
  2. (2)
    The partial bijection $\alpha_{s}$ is a *-homomorphism.
Given an idempotent $e\in E_{\mathcal{S}}$, we set $I_{e}=\text{dom}(\alpha_{e})$. We observe that $\alpha_{s}$ is a *-isomorphism of $I_{s^{*}s}$ onto $I_{ss^{*}}$. The next example, central in [lledomartinez], is also the main C-action we will consider in this note.
Example 2.1.
Let $\mathcal{A}=\ell^{\infty}(\mathcal{S})$. For each $e\in E_{\mathcal{S}}$, define $I_{e}=\{f\in\mathcal{A}:\text{supp}(f)\subseteq e\mathcal{S}\}$. For each $s\in\mathcal{S}$, define $\alpha_{s}:I_{s^{*}s}\to I_{ss^{*}}$ by $\alpha_{s}f](t)=f(s^{*}t)$. The mapping $s\in\mathcal{S}\mapsto\alpha_{s}\in\mathcal{I}(\mathcal{S})$ is an action of $\mathcal{S}$ on $\mathcal{A}$.
The C-algebra $\ell^{\infty}(\mathcal{S})$ is *-isomorphic to $C(\beta\mathcal{S})$, where $\beta\mathcal{S}$ is the Stone-Čech compactification of $\mathcal{S}$. At the level of compact Hausdorff spaces, inverse semigroup actions are described as follows (see [EXEL] for details): an action of $\mathcal{S}$ on a compact Hausdorff space $\mathcal{X}$ is a triple $(\mathcal{S},\mathcal{X},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{X})$ is a semigroup homomorphism such that, for all $s\in\mathcal{S}$,
@@
[…]
Example 2.2 ([gondek24], Section 3.2).
@@
Let us define an inverse semigroup action of $\mathcal{S}$ on the Stone-Čech compactification $(\beta\mathcal{S},\iota)$, where $\mathcal{S}$ is viewed as a discrete space and $\iota:\mathcal{S}\to\beta\mathcal{S}$ is the canonical embedding. For each $s\in\mathcal{S}$, define a bijection $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ by $\lambda_{s}t=st$ ($t\in s^{*}\mathcal{S}$). By the universal property of $\beta\mathcal{S}$, $\lambda_{s}$ extends to a homeomorphism $\lambda_{s}:\overline{s^{*}\mathcal{S}}\to\overline{s\mathcal{S}}$, where $\overline{s^{*}\mathcal{S}}:=\overline{\iota(s^{*}\mathcal{S})}^{\beta}\subseteq\beta\mathcal{S}$. We observe that the domain and codomain of $\lambda_{s}$ are always a compact, open subsets of $\beta\mathcal{S}$. It is routine to check that the mapping $\mathcal{S}\ni s\mapsto\lambda_{s}\in\mathcal{I}(\beta\mathcal{S})$ is a semigroup homomorphism, and therefore an action, which we shall henceforth refer to by $\mathcal{S}\curvearrowright\beta\mathcal{S}$.
The contravariant action $\hat{\lambda}$ of $\mathcal{S}$ on $C(\beta\mathcal{S})$ is the action in Example 2.1, after applying the identification $j:C(\beta\mathcal{S})\ni f\mapsto f\circ\iota\in\ell^{\infty}(\mathcal{S})$. Indeed, given $f\in I_{s^{*}s}\subseteq\ell^{\infty}(\mathcal{S})$, we have $j(f)](t)=f(\iota(t))=0\Leftrightarrow\iota(t)\not\in\overline{s^{*}\mathcal{S}}\Leftrightarrow t\not\in s^{*}\mathcal{S}$, so that $j(\text{dom}(\hat{\lambda}_{s}))=I_{s^{*}s}$, and
$[\alpha_{s}(j(f))](t)=[j(f)](s^{*}t)=f(\iota(s^{*}t))=f(\lambda_{s^{*}}\iota(t))=[\hat{\lambda}_{s}f](t),$
for all $f\in I_{s^{*}s}$.
Given a C-algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ and an inverse semigroup action $\mathcal{S}\curvearrowright\mathcal{A}$, there is a standard construction for a reduced crossed product C-algebra $\mathcal{A}\rtimes_{r}\mathcal{S}\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$, which does not depend on the inclusion $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ (for details, see [EXEL]). We will not work directly with this C-algebra, so we only sketch the construction. For each $s\in\mathcal{S}$, we define the maps $V_{s}:\mathcal{S}\to\mathcal{B}(\ell^{2}(\mathcal{S}))$ and $\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}\otimes\mathcal{\ell}^{2}(\mathcal{S}))$ by
$V_{s}\delta_{t}=\begin{cases}\delta_{st}&t\in s^{*}\mathcal{S}\\ 0&\text{otherwise}\end{cases}\qquad$ $\text{and}\qquad\pi(a)(\xi\otimes\delta_{t})=\begin{cases}\alpha_{t^{*}}(a)\xi\otimes\delta_{t}&t\in I_{s^{*}s}\\ 0&\text{otherwise}.\end{cases}$
The reduced crossed product is the C-subalgebra of $\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$ defined by
$\mathcal{A}\rtimes_{r}\mathcal{S}:=\textup{C}^{*}(\{1\otimes V_{s}:s\in\mathcal{S}\}\cup\pi(A))\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S})).$
@@
We are interested in describing an étale groupoid $\mathcal{G}$ with the property that $\text{C}^{*}_{r}(\mathcal{G})=\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. In preparation for this, let us establish conventions for étale groupoids in this note. Given a groupoid $\mathcal{G}$, we will denote by $\mathcal{G}^{(2)}\subset\mathcal{G}\times\mathcal{G}$ the composable pairs, and $\mathcal{G}^{(0)}=\{\gamma^{-1}\gamma:\gamma\in\mathcal{G}\}$ the unit space. The source and range maps $\mathbf{d}:\mathcal{G}\to\mathcal{G}^{(0)}$ and $\mathbf{r}:\mathcal{G}\to\mathcal{G}^{(0)}$ (respectively) are defined by $\mathbf{d}(\gamma)=\gamma^{-1}\gamma$ and $\mathbf{r}(\gamma)=\gamma\gamma^{-1}$, for $\gamma\in\mathcal{G}$. Given $u\in\mathcal{G}^{(0)}$, we will write $\mathcal{G}_{u}=\mathbf{d}^{-1}(u)$, $\mathcal{G}^{u}=\mathbf{r}^{-1}(u)$, and $\mathcal{G}^{u}_{u}=\mathcal{G}_{u}\cap\mathcal{G}^{u}$. The isotropy $\textup{Iso}(\mathcal{G})$ is defined by
@@
A topological groupoid is a groupoid $\mathcal{G}$ equipped with a topology under which the multiplication $\mathcal{G}^{(2)}\ni(\alpha,\beta)\mapsto\alpha\beta$ and inverse map $\mathcal{G}\ni\gamma\mapsto\gamma^{-1}$ are continuous. A topological groupoid is called étale if it is locally compact, $\mathcal{G}^{(0)}$ is Hausdorff in the subspace topology inherited from $\mathcal{G}$, and the maps $\mathbf{d}$ and $\mathbf{r}$ are local homeomorphisms. We remind the reader that étale groupoids are not required to be Hausdorff. It is a standard fact that an étale groupoid $\mathcal{G}$ is Hausdorff if and only if $\mathcal{G}^{(0)}$ is closed.
Now let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a locally compact Hausdorff space. We associate an étale groupoid to $(\mathcal{S},\mathcal{X},\alpha)$ as follows (see [EXEL] for details). Set
@@
Every étale groupoid $\mathcal{G}$ generates a reduced C-algebra, $\textup{C}^{*}_{r}(\mathcal{G})$. As we will not have need to directly work with this C-algebra, we refer the reader to [CEPSS] for details. The relation between the two C-algebra constructions described in this section is the content of the next theorem, which is an immediate consequence of ([bussexel], Theorem 4.11).
Theorem 2.3.
Let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a compact Hausdorff space. The reduced crossed product $C_{0}(\mathcal{X})\rtimes_{r}\mathcal{S}$ is isomorphic to the reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes_{\alpha}\mathcal{X})$.

### v2

@@ Abstract.
We initiate the study of the Stone-Čech transformation groupoid $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ of an inverse semigroup $\mathcal{S}$. We prove that the properties of being Hausdorff, principal, and effective are all equivalent for $\mathcal{G}$, and give an algebraic condition on $\mathcal{S}$ equivalent to the Hausdorffness of $\mathcal{G}$. We show that the Hausdorffness of Exel’s tight groupoid $\mathcal{G}_{\text{tight}}(\mathcal{S})$ is necessary for the Hausdorffness of $\mathcal{G}$. Finally, we clarify the connection between several crossed product constructions involving this groupoid, and show that when $\mathcal{G}$ is Hausdorff and $\mathcal{S}$ has the Property (FL) of Lledó and Martínez, then $\mathcal{G}$ is amenable if and only if the reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
@@ 1. Introduction
A semigroup $\mathcal{S}$ is called inverse if, for every $s\in\mathcal{S}$, there is a unique $s^{*}\in\mathcal{S}$ with $ss^{*}s=s$ and $s^{*}ss^{*}=s^{*}$. There have been recent efforts to understand the dynamics and representation theory of inverse semigroups in a way that generalizes the discrete group setting. The first to investigate the exactness of $\textup{C}^{*}_{r}(\mathcal{S})$ seems to be Anantharaman-Delaroche in the monograph [2], where groupoid C-algebra techniques are applied to show that $\textup{C}^{*}_{r}(\mathcal{S})\subset\mathcal{B}(\ell^{2}(\mathcal{S}))$ is exact whenever the maximum group homomorphic image $\mathcal{S}{/}\sigma$ is an exact group. In [14], Lledó and Martínez construct and study the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}\subset\mathcal{B}(\ell^{2}(\mathcal{S}))$ of a countable inverse semigroup $\mathcal{S}$. Under a mild combinatorial assumption (admitting a finite labelling), they show that the reduced inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{R}_{\mathcal{S}}$ is nuclear.
In this note, we consider the contravariant action $\mathcal{S}\curvearrowright\beta\mathcal{S}$, and investigate the properties of the resulting action groupoid $\mathcal{S}\ltimes\mathcal{\beta}\mathcal{S}$ and reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes\mathcal{\beta}\mathcal{S})\cong\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. We first consider the question of effectiveness and principality for $\mathcal{S}\ltimes\beta\mathcal{S}$. These conditions always hold in the group case, and correspond to the fact that the action $\Gamma\curvearrowright\beta\Gamma$ is always free (see for example [15]*Example 2.14). In contrast to the group case, and somewhat surprisingly, there are inverse semigroups $\mathcal{S}$ such that $\mathcal{S}\ltimes\beta\mathcal{S}$ is not effective nor principal. Using a finite cover criterion of Exel and Pardo [7], we obtain the following generalization of the group case:
@@
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
@@
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is effective;
  3. (3)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal;
@@
This generalizes the group case because $\mathcal{S}\ltimes\beta\mathcal{S}$ is always Hausdorff when $\mathcal{S}$ is a group. Using Theorem 1, we give an algebraic criterion on $\mathcal{S}$ (Theorem 3.7) which implies $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff for the class of abstract pseudogroups (Corollary 3.8). Finally, since there are many examples of inverse semigroups with non-Hausdorff tight groupoids, condition (3) of Theorem 1 provides many examples of inverse semigroups for which $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff nor principal.
To consider the amenability of the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ in the sense of Exel and Starling [8], we clarify the relationship between the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$ and the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ in the sense of [4]. In [14], the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}\subset\mathcal{B}(\ell^{2}(\mathcal{S}))$ is shown to embed into $\mathcal{B}(\ell^{2}(\mathcal{S})\otimes\ell^{2}(\mathcal{S}))$ as the image of a covariant representation of the action $\mathcal{S}\curvearrowright\ell^{\infty}(\mathcal{S})$. It follows from Theorem 1 that this C-algebra, denoted in [14] by $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$” and in this note by $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$, is not to be confused (in full generality) with the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ in the sense of [4]. However, these C-algebras often coincide:
Theorem 2.
The following conditions are equivalent for an inverse semigroup $\mathcal{S}$:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}\cong\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ canonically.
As a consequence, we obtain a generalization of a discrete group result.
Theorem 3.
Suppose $\mathcal{S}$ is a countable inverse semigroup with the property (FL) of Lledó and Martínez. If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable (in the sense of [8]) if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
We would like to thank the anonymous referee for several suggestions and corrections from the previous version to this manuscript, and particularly for pointing out how our results in that version suggest the content of Theorem 2.
@@ 2. Preliminaries
In what follows, $\mathcal{S}$ denotes a countable inverse semigroup. We refer the reader to Lawson’s book [13] for a detailed treatment of these semigroups. A subset $A\subset\mathcal{S}$ is called symmetric if $s^{*}\in A$ whenever $s\in A$. An element $e\in\mathcal{S}$ is called idempotent if $e^{2}=e$. For any $\in\mathcal{S}, ss$$^{*}s$ and $ss^{*}$ are idempotent. Denote by $E_{\mathcal{S}}$ the set of idempotent elements of $\mathcal{S}$. It is a pleasant fact that $E_{\mathcal{S}}$ is a commutative subsemigroup of $\mathcal{S}$, and that $\mathcal{S}$ is a group if and only if $E_{\mathcal{S}}$ is a singleton. If $X$ is a set, then the collection $\mathcal{I}(X)$ of all partial bijections between subsets of $X$ is an inverse semigroup under the operation of partial composition. The semigroup $\mathcal{I}(X)$ is called the symmetric inverse monoid on $X$.
@@
For any poset $(P,\prec)$ and subset $A\subseteq P$ we will use the notation
@@
An action of $\mathcal{S}$ on a compact Hausdorff space $\mathcal{X}$ is a triple $(\mathcal{S},\mathcal{X},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{X})$ is a semigroup homomorphism such that, for all $s\in\mathcal{S}$,
@@
One can see that the domain of $\alpha_{s}$ is the same as that of $\alpha_{s^{*}s}$. For each $e\in E_{\mathcal{S}}$ we write $D_{e}^{\alpha}$ for the domain of $\alpha_{e}$ (or simply $D_{e}$ if the action is understood), so that $\alpha_{s}:D_{s^{*}s}^{\alpha}\to D_{ss^{*}}^{\alpha}$. The main action we are concerned with in this paper is described below.
Example 2.1 ([11], Section 3.2).
@@
We define the canonical action of an inverse semigroup $\mathcal{S}$ on its Stone-Čech compactification $(\beta\mathcal{S},\iota)$, where $\mathcal{S}$ is viewed as a discrete space and $\iota:\mathcal{S}\to\beta\mathcal{S}$ is the canonical embedding. For each $s\in\mathcal{S}$, define a bijection $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ by $\lambda_{s}t=st$ ($t\in s^{*}\mathcal{S}$). By the universal property of $\beta\mathcal{S}$, $\lambda_{s}$ extends to a homeomorphism $\lambda_{s}:\overline{s^{*}\mathcal{S}}\to\overline{s\mathcal{S}}$, where $\overline{s^{*}\mathcal{S}}:=\overline{\iota(s^{*}\mathcal{S})}^{\beta}\subseteq\beta\mathcal{S}$. We observe that the domain and codomain of $\lambda_{s}$ are always compact, open subsets of $\beta\mathcal{S}$. It is routine to check that the mapping $\mathcal{S}\ni s\mapsto\lambda_{s}\in\mathcal{I}(\beta\mathcal{S})$ is a semigroup homomorphism, and therefore an action, which we shall henceforth refer to by $\mathcal{S}\curvearrowright\beta\mathcal{S}$.
@@
The appropriate way to interpret the dynamics of an inverse semigroup action $\mathcal{S}\curvearrowright\mathcal{X}$ is through the corresponding transformation groupoid, denoted by $\mathcal{S}\ltimes\mathcal{X}$. Before reminding the reader of this construction, we establish our conventions for étale groupoids. We direct the reader to the lecture notes [21] of Sims for details. A groupoid is a small category in which every arrow is invertible. Given a groupoid $\mathcal{G}$, we will denote by $\mathcal{G}^{(2)}\subset\mathcal{G}\times\mathcal{G}$ the composable pairs, and $\mathcal{G}^{(0)}=\{\gamma^{-1}\gamma:\gamma\in\mathcal{G}\}$ the unit space. The source and range maps $\mathbf{d}:\mathcal{G}\to\mathcal{G}^{(0)}$ and $\mathbf{r}:\mathcal{G}\to\mathcal{G}^{(0)}$ (respectively) are defined by $\mathbf{d}(\gamma)=\gamma^{-1}\gamma$ and $\mathbf{r}(\gamma)=\gamma\gamma^{-1}$, for $\gamma\in\mathcal{G}$. Given $u\in\mathcal{G}^{(0)}$, we will write $\mathcal{G}_{u}=\mathbf{d}^{-1}(u)$, $\mathcal{G}^{u}=\mathbf{r}^{-1}(u)$, and $\mathcal{G}^{u}_{u}=\mathcal{G}_{u}\cap\mathcal{G}^{u}$. The isotropy $\textup{Iso}(\mathcal{G})$ is defined by
@@
A topological groupoid is a groupoid $\mathcal{G}$ equipped with a topology under which the multiplication $\mathcal{G}^{(2)}\ni(\alpha,\beta)\mapsto\alpha\beta$ and inverse map $\mathcal{G}\ni\gamma\mapsto\gamma^{-1}$ are continuous. A topological groupoid is called étale if it is locally compact, $\mathcal{G}^{(0)}$ is Hausdorff in the subspace topology inherited from $\mathcal{G}$, and the maps $\mathbf{d}$ and $\mathbf{r}$ are local homeomorphisms. A bisection is a subset $U\subset\mathcal{G}$ such that $\mathbf{d}|_{U}$ and $\mathbf{r}|_{U}$ are local homeomorphisms. We remind the reader that étale groupoids are not required to be Hausdorff. It is a standard fact that an étale groupoid $\mathcal{G}$ is Hausdorff if and only if $\mathcal{G}^{(0)}$ is closed.
Now let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a locally compact Hausdorff space. We associate an étale groupoid to $(\mathcal{S},\mathcal{X},\alpha)$ as follows (see [9] for details). Set